Grade 12th passInorganic Chemistry

Why is cobalts oxidation potential from Co2+/Co3+ is very high....

VIDYARANI SINGADI
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago

[Co(H2O)6]2+ is more stable than [Co(H2O)6]3+ because Co2+ is more stabilised by weak field ligand, H2O than is Co3+.

Co2+(3d7)= (t2g)5(eg)2 has higher value of CFSE than Co3+(d6) = (t2g)4(eg)2 in the weak octahedral field leading to greater stability of Co2+(aq) than Co2+(aq).

On the other hand, Co(NH3)6]3+ is more stable than [Co(NH3)6]2+ because Co3+ is more stabilised by strong field ligand NH3 than is Co2+.

This is due to higher value of CFSE for Co3+ (3d6) = (t2g)6(eg)0 than Co2+ (3d7) = (t2g)6(eg)1 in the strong octahedral field ligand NH3 leading to greater stability in former.

The calculations are done on the basis of CFT.